# Conventional Quantum Theory Does Not Support A Coherent Relational Account

### v1

@@ 2. Critique
Now, let $W$ subject the $B+A$ system to a measurement of the Bell observable for which the state $|\Phi^{+}\rangle$ is an eigenstate. The Bell observable measurement is somewhat involved, but there is nothing to preclude it in principle. $W$ could comprise as many bilevel systems as one wishes in order to allow for a pointer, signaling system, and any other degrees of freedom needed to implement the required correlations via interaction Hamiltonians, under the assumption (UOA) that all that is required for a “measurement outcome” (or outcomes in an Everettian approach) is a one-to-one correlation between eigenstates of an observable of the system being measured and the states of the putative “observer” (as assumed in Proietti et al). [^8] The Bell observable is not uniquely required. Alternatively, one could arrange for $F$’s pointer $B$ to be anti-correlated with $A$’s putative outcomes, and have $W$ conduct a measurement of the total spin of the system. This would reveal the inconsistency since $W$ predicts that the total spin $J$ must always be found to be 1, while according to F, W has a 50% chance of finding the result $J=0$, for the singlet state. $W$’s measurement is accompanied by a signal to $F$ as follows. An outcome finding the state $|\Phi^{+}\rangle$ results in a photon being emitted to $F$ to excite his communication degree of freedom $C$. This could be implemented by having $W$’s outcome registration involve excitation of a degree of freedom with a very short decay time such that it immediately de-excites in a collimated direction to $F$. According to $W$, $F$ should receive that photon for every run of the experiment. This makes the inconsistency manifest, since according to F, his probability of receiving the photon is only 1/2 given his description of the situation by the mixed state (1,2). (This concludes material excerpted from Kastner, 2022.)
We should note that the ”basic postulate of RQM” (Rovelli 1986), repeated by Laudisa and Rovelli 2021) is refuted by this counterexample:
@@
Thus, RQM asserts that $F$ should be able to predict correctly the future values of its own local system(s) by reference only to its local past outcomes, while its presumed unitary entanglement with the result of the world continues. But in the counterexample (presuming that $W$ is correct, as AR assert), $F$ cannot correctly predict the probability of a future event relative to itself—i.e., the probability of its degree of freedom $C$ becoming excited—based on its own past outcomes. This probability apparently depends instead on another system outside ($W$), contrary to the central postulate of RQM. At this point, RQM is effectively dead. The version offered in AR2022, RQM22, is a mere shadow of its former self. Yet even that shadow fails to survive, as we shall see in what follows.
@@
“Now suppose Bob measures S in the same basis as Alice’s measurement, and hence he obtains a measurement outcome MBS which he will interpret as providing information about the result of Alice’s measurement on S. Suppose that Bob also ‘measures’ Alice herself and obtains a measurement outcome MBA for the value of of some pointer variable which is supposed to be a record of her measurement result - for example, he could simply ask her what her measurement result was. So in this scenario we have three different measurement outcomes MA,MBS,MBA all supposedly providing information about the value of the same variable.” (AR2022)
@@
Definition 4.1. Cross-perspective links: In a scenario where some observer Alice measures a variable V of a system S, then provided that Alice does not undergo any interactions which destroy the information about V stored in Alice’s physical variables, if Bob subsequently measures the physical variable representing Alice’s information about the variable V, then Bob’s measurement result will match Alice’s measurement result.” (AR20220
The provision about destroyed information does a lot of ‘heavy lifting’ in the authors’ subsequent attempt to remove the threat to RQM of the Wigner’s Friend paradox (as we will discuss further below). However, the system in the counterexample has a degree of freedom C that is not subject to any interaction that would ‘destroy’ its information, and F makes a specific prediction about the probability of a state change in C that disagrees with the probability predicted by W. If one takes the ‘correct’ description as the that of F being in a macroscopic superposition, then F must find out he is wrong, violating the central postulate of RQM as noted above. Does RQM22 do any better? AR2022 offer the following:
@@
However, in the above counterexample, according to CQM (which assumes there is no real reduction), the inner observer will indeed find that he is ‘wrong’ just by checking his degree of freedom C. There is no ‘box’ to step out of, nor any mechanism for wiping his memory. The assumption that the inner observer’s memory will be wiped is merely an article of faith that is falsified by the counterexample. RQM22 fares no better than RQM, falling victim to CQM’s acknowledged inability to account for determinate outcomes.
@@
As noted above, the authors assert that Wigner’s outcome is an ‘observer-independent fact’ (BF sentence). Yet CQM fails to license this assumption, since one could always have another ‘outside observer’ Wigner2 who (according to CQM) must describe Wigner as being in a coherent superposition after his measurement—since according to CQM, no observer is ever correct in thinking that any collapse to an eigenvalue actually occurred. This means that Wigner’s outcome cannot be an ‘observer independent fact’ (at least as a piece of information supporting a correct prediction) as asserted in the quoted excerpt. This point reveals the fallacy in the following argumentation by the authors:
“Note also that Bob’s erroneous prediction arises only because he assumes that there is some kind of collapse when he performs his measurement. If instead he uses the version of RQM that we have suggested in this article, then he will correctly predict that Wigner could still get the result 1, since he will be aware that although his observation has caused him to update his own relative state, he remains in a superposition state relative to Wigner (since that state describes the relational [sic] between Wigner and Bob, not the absolute state of Bob). So in fact no contradictory predictions will arise as long as Bob and Wigner are consistent about employing the same interpretation of quantum mechanics.” (AR2022)
The infinite regress becomes obvious here if we simply replace Bob by Wigner and Wigner by Wigner2:
@@
In other words, the ostensible remedy for the MP-related inconsistencies, according to RQM, is to assume that every observer’s prediction, based on his measurement outcome and assignment of the corresponding eigenstate, is potentially erroneous based on the possible existence of an ‘outer observer.’ Along with this goes the untenable assumption that the inner observer will have no memory of being wrong, which is falsified by the counterexample discussed above. Thus, RQM prescribes that all experimenters are wrong when they attribute outcome eigenstates to their systems, since some outside observer would deny that attribution. But there could never be any correct ‘outside observer’ due to the infinite regress of ‘outside observers’. And indeed there can be no ‘observer independent facts’ if all observers are mistaken whenever they think that they got an outcome corresponding to an eigenstate of the measured observable. RQM thus fails to support any observer-independent facts, contrary to the authors’ claim, i.e.:
“…we now regard the pointlike quantum events or ‘flashes’ as absolute, observer-independent facts about reality, rather than relativizing them to an observer.” (AR2022)
This simply will not do, since there is no physics given to support such a ‘quantum event’. If Bob makes a measurement and is taken as finding an outcome, the authors seem to want to take that as a ‘quantum event’ corresponding to a ‘fact’. But under CQM (and RQM and RQM22), that ‘fact’ must be wrong, and so is Wigner’s, ad infinitum. In what coherent sense could myriads of wrong impressions possibly count as ‘observer-independent facts about reality’?
A further note is in order regarding the approach advocated in Baumann and Brukner (2019) (henceforth BB2019), which seeks to tame the WF-type inconsistencies by recourse to an instrumentalist treatment. They consider essentially the same counterexample discussed in Section 2b, in which an ‘inner observer’ could (according to CQM) find out they were wrong about their outcome-based prediction. BB2019 proceed under the conventional assumption that CQM is the correct theory and that the resulting WF-type scenarios are arguments against the existence of ‘observer-independent facts’, as in their remark:
@@
They then offer an instrumentalist approach involving ‘rational agents’ to attempt to accommodate the resulting inconsistencies, in which agents can update their ‘degree of belief’ subject to new information from an ‘outside observer’. However, BB2019 allow for an ‘automatic machine’ to serve as ‘Wigner’ and to obtain an outcome which it then passes to the Friend (to get around worries about Wigner being dishonest about his outcome). The authors appear not to notice that they have just helped themselves to an observer-independent fact, namely the machine’s outcome. In other words, in attributing an outcome to a machine, they tacitly assume that quantum theory does in fact apply to physical systems, not ‘agents’, so that a ‘fact’ can obtain without an agent. Which is it? Are the ‘facts’ corresponding to outcomes observer-dependent or not? The discussion in BB2019 assumes both, despite the fact that they are mutually exclusive. Moreover, in any case, CQM fails to predict any outcomes. If one helps oneself to an outcome despite that, the outcome (or at least any ‘fact’ it might be taken to imply) is subject to being wrong, whether it is attributed to an agent or to a physical system. And the latter has no ‘degree of belief’ to update. The theory—CQM—that is unaccountably taken as yielding an outcome (or ‘fact’) simply cannot provide a correct one.
@@ 3. Conclusion
The above points should also be kept in mind when evaluating the preferred options of Baumann and Brukner (2019), who treat quantum theory in an instrumentalist fashion, as a tool for making predictions. They refer to the inner observer as being subject to ‘updating degrees of belief’, but as noted above, under that approach there is still in principle an infinite regress of observers whose ‘observed outcomes’ are neither predicted by CQM nor can be regarded as facts about reality. Yet even the latter treatment implicitly smuggles in the observer-independent occurrence of outcomes in assuming that an ‘automatic machine’ could play the part of the outer observer. In that case, the infinite regress applies, and the ‘automatic machine’ could somehow be ‘mistaken’ about its outcome. Thus, once again, CQM is faced with a fatal inconsistency among outcomes that no resort to ‘updating’ can cure, since a machine is not subject to ‘degrees of belief’.
In contrast to this unsatisfactory situation facing conventional quantum theory and its putative interpretation RQM, it has been recalled that theories that incorporate objective reduction are not subject to the measurement problem (as acknowledged by one of the authors in a separate publication, Adlam 2023) and thus are not subject to its attendant paradoxes that lead to inconsistencies subject to remedy only by resort to observer-dependent instrumentalism and ad hoc postulations of outcomes not predicted by the theory itself, and in which all such postulated outcomes are subject to yielding wrong predictions.

### v2

@@ 2. Critique
Now, let $W$ subject the $B+A$ system to a measurement of the Bell observable for which the state $|\Phi^{+}\rangle$ is an eigenstate, where only the outcome $\Phi^{+}$ triggers the emission of a photon to $F$ to excite his communication degree of freedom $C$ [^9] The Bell observable is not uniquely required. Alternatively, one could arrange for $F$’s pointer $B$ to be anti-correlated with $A$’s putative outcomes, and have $W$ conduct a measurement of the total spin of the system. This would reveal the inconsistency since $W$ predicts that the total spin $J$ must always be found to be 1, while according to F, W has a 50% chance of finding the result $J=0$, for the singlet state. According to $W$, $F$ should receive that photon for every run of the experiment. However, according to $F$, his probability of receiving the photon is only 1/2 in view of his description of the situation by the mixed state (1,2). This makes the inconsistency manifest.
We should note that the ”basic postulate of RQM” (Rovelli 1986), repeated by Laudisa and Rovelli 2021) is refuted by this example:
@@
Thus, RQM asserts that $F$ should be able to predict correctly the future values of its own local system(s) by reference only to its local past outcomes, while its presumed unitary entanglement with the result of the world continues. But in the counterexample (presuming that $W$ is correct, as Adlam and Rovelli assert), $F$ cannot correctly predict the probability of a future event relative to itself—i.e., the probability of its degree of freedom $C$ becoming excited—based on its own past outcomes. This probability apparently depends instead on another system outside ($W$), contrary to the central postulate of RQM. At this point, RQM is effectively dead. The version offered in Adlam and Rovelli 2022, RQM22, is a mere shadow of its former self. Yet even that shadow fails to survive, as we shall see in what follows.
@@
“Now suppose Bob measures S in the same basis as Alice’s measurement, and hence he obtains a measurement outcome MBS which he will interpret as providing information about the result of Alice’s measurement on S. Suppose that Bob also ‘measures’ Alice herself and obtains a measurement outcome MBA for the value of of some pointer variable which is supposed to be a record of her measurement result - for example, he could simply ask her what her measurement result was. So in this scenario we have three different measurement outcomes MA,MBS,MBA all supposedly providing information about the value of the same variable.” (Adlam and Rovelli 2022)
@@
Definition 4.1. Cross-perspective links: In a scenario where some observer Alice measures a variable V of a system S, then provided that Alice does not undergo any interactions which destroy the information about V stored in Alice’s physical variables, if Bob subsequently measures the physical variable representing Alice’s information about the variable V, then Bob’s measurement result will match Alice’s measurement result.” (Adlam and Rovelli 2022
The provision about destroyed information does a lot of ‘heavy lifting’ in the authors’ subsequent attempt to remove the threat to RQM of the Wigner’s Friend paradox (as we will discuss further below). However, the system in the counterexample has a degree of freedom C that is not subject to any interaction that would ‘destroy’ its information, and F makes a specific prediction about the probability of a state change in C that disagrees with the probability predicted by W. If one takes the ‘correct’ description as that of F being in a pure state superposition, then F must find out he is wrong, violating the central postulate of RQM as noted above. Does RQM22 do any better? The authors offer the following:
@@
However, in the above counterexample, according to Unitary Quantum Theory, the inner observer will indeed find that he is ‘wrong’ just by checking his degree of freedom C. There is no ‘box’ to step out of, nor any mechanism for wiping his memory. The assumption that the inner observer’s memory will be wiped is merely an article of faith that is falsified by the counterexample. RQM22 fares no better than RQM, falling victim to the acknowledged inability of Unitary Quantum Theory to account for determinate outcomes.
@@
As noted above, the authors assert that Wigner’s outcome is an ‘observer-independent fact’ (boldface sentence). Yet Unitary Quantum Theory fails to license this assumption, since one could always have another ‘outside observer’ Wigner2 who (according to Unitary Quantum Theory) must describe Wigner as being in a coherent superposition after his measurement—since according to Unitary Quantum Theory, no observer is ever correct in thinking that any collapse to an eigenvalue actually occurred. This means that Wigner’s outcome cannot be an ‘observer independent fact’ (at least as a piece of information supporting a correct prediction) as asserted in the quoted excerpt. This point reveals the fallacy in the following argumentation by the authors:
“Note also that Bob’s erroneous prediction arises only because he assumes that there is some kind of collapse when he performs his measurement. If instead he uses the version of RQM that we have suggested in this article, then he will correctly predict that Wigner could still get the result 1, since he will be aware that although his observation has caused him to update his own relative state, he remains in a superposition state relative to Wigner (since that state describes the relational [sic] between Wigner and Bob, not the absolute state of Bob). So in fact no contradictory predictions will arise as long as Bob and Wigner are consistent about employing the same interpretation of quantum mechanics.” (Adlam and Rovelli 2022)
The infinite regress becomes obvious here if we simply replace Bob by Wigner and Wigner by Wigner2, who measures some other observable $X$:
@@
In other words, the ostensible remedy for the inconsistencies, according to RQM, is to assume that every observer’s prediction, based on his putative measurement outcome and assignment of the corresponding eigenstate, is potentially erroneous based on the possible existence of an ‘outer observer.’ Along with this goes the untenable assumption that the inner observer will have no memory of being wrong, which is falsified by the counterexample discussed above. Thus, RQM prescribes that all experimenters are wrong when they attribute outcome eigenstates to their systems, since some outside observer would deny that attribution. But there could never be any correct ‘outside observer’ due to the infinite regress of ‘outside observers’. And indeed there can be no ‘observer independent facts’ if all observers are mistaken whenever they think that they got an outcome corresponding to an eigenstate of the measured observable. RQM thus fails to support any observer-independent facts, contrary to the authors’ claim, i.e.:
“…we now regard the pointlike quantum events or ‘flashes’ as absolute, observer-independent facts about reality, rather than relativizing them to an observer.” (Adlam and Rovelli 2022)
This simply will not do, since there is no physics given to support such a ‘quantum event’. If Bob makes a measurement and is taken as finding an outcome, the authors seem to want to take that as a ‘quantum event’ corresponding to a ‘fact’. But under Unitary Quantum Theory (and RQM and RQM22), that ‘fact’ must be wrong, and so is Wigner’s, ad infinitum. In what coherent sense could myriads of wrong impressions possibly count as ‘observer-independent facts about reality’?
A further note is in order regarding the approach advocated in Baumann and Brukner (2019) (henceforth BB2019), which seeks to tame the Wigner’s Friend-type inconsistencies by recourse to an instrumentalist treatment. They consider essentially the same counterexample discussed in Section 2b, in which an ‘inner observer’ could (according to Unitary Quantum Theory) find out they were wrong about their outcome-based prediction. BB2019 proceed under the conventional assumption that Unitary Quantum Theory is the correct theory and that the resulting Wigner’s Friend-type scenarios are arguments against the existence of ‘observer-independent facts’, as in their remark:
@@
They then offer an instrumentalist approach involving ‘rational agents’ to attempt to accommodate the resulting inconsistencies, in which agents can update their ‘degree of belief’ subject to new information from an ‘outside observer’. However, BB2019 allow for an ‘automatic machine’ to serve as ‘Wigner’ and to obtain an outcome which it then passes to the Friend (to get around worries about Wigner being dishonest about his outcome). The authors appear not to notice that they have just helped themselves to an observer-independent fact, namely the machine’s outcome. In other words, in attributing an outcome to a machine, they tacitly assume that quantum theory does in fact apply to physical systems, not ‘agents’, so that a ‘fact’ can obtain without an agent. Which is it? Are the ‘facts’ corresponding to outcomes observer-dependent or not? The discussion in BB2019 assumes both, despite the fact that they are mutually exclusive. Moreover, in any case, Unitary Quantum Theory fails to predict any outcomes. If one helps oneself to an outcome despite that, the outcome (or at least any ‘fact’ it might be taken to imply) is subject to being wrong, whether it is attributed to an agent or to a physical system. And the latter has no ‘degree of belief’ to update. The theory—Unitary Quantum Theory—that is unaccountably taken as yielding an outcome (or ‘fact’) simply cannot provide a correct one.
@@ 3. Conclusion
The above points should also be kept in mind when evaluating the preferred options of Baumann and Brukner (2019), who treat quantum theory in an instrumentalist fashion, as a tool for making predictions. They refer to the inner observer as being subject to ‘updating degrees of belief’, but as noted above, under that approach there is still in principle an infinite regress of observers whose ‘observed outcomes’ are neither predicted by Unitary Quantum Theory nor can be regarded as facts about reality. Moreover, the latter treatment implicitly smuggles in the observer-independent occurrence of outcomes in assuming that an ‘automatic machine’ could play the part of the outer observer. Yet in that case, the infinite regress still applies, and the ‘automatic machine’ could somehow be ‘mistaken’ about its outcome. Thus, once again, Unitary Quantum Theory is faced with a fatal inconsistency among outcomes that no resort to ‘updating’ can cure, since a machine is not subject to ‘degrees of belief’.
In contrast to this unsatisfactory situation facing conventional unitary quantum theory and its putative interpretation RQM, it has been recalled that theories that incorporate objective reduction are not subject to the measurement problem (as acknowledged by one of the authors in a separate publication, Adlam 2023) and thus are not subject to its attendant paradoxes that lead to inconsistencies subject to remedy only by resort to observer-dependent instrumentalism and ad hoc postulations of outcomes not predicted by the theory itself, and in which all such postulated outcomes are subject to being incorrect.
Acknowledgments. I am grateful to two anonymous referees for valuable suggestions for improvement of the presentation.